OpenStudy (anonymous):

What is the surface area of a circle that has a diameter of 14 inches?

OpenStudy (cggurumanjunath):

http://www.math.com/tables/geometry/surfareas.htm

Directrix (directrix):

@MathFrantic A circle has area but not surface area. Surface area refers to 3-D solids such as a sphere. The diameter of a circle is twice the radius of the circle. So, if the diameter is 14, then the radius of the circle is 7. The area formula for a circle is the following: A = π * r² where r is the radius of the circle. A = π * 7² So crank that out for the area. Post what you get and we can compare answers. Okay?

OpenStudy (anonymous):

So if the radius is 7 you square that which gives you 49, then multiply by pie??

Directrix (directrix):

Correct. I don't know if your answer options are in exact form or are approximated.

OpenStudy (anonymous):

153.86 in2 is one of the choices, but the in2 is whats confusing me..

Directrix (directrix):

Area is expressed in square inches. A = pi * 7 inches * 7 inches A = 49 pi * inches² Inches times inches is square inches just as 7 * 7 = 7²

Directrix (directrix):

I got 153.94 square inches but I may have used more digits of pi than you.

OpenStudy (anonymous):

Okay, Thank you for the help!